OpenStudy (anonymous):

There are 5 green balls 1 yellow ball and 4 red balls in a bag what are the following probabilities Choosing red: 2/6? Not choosing yellow: 9\10 choosing blue: ?

OpenStudy (anonymous):

there were no blue balls mentioned so 0

OpenStudy (anonymous):

Thanks is everything else good

OpenStudy (anonymous):

Choosing red= 4/10=2/5. not 2/6.

OpenStudy (anonymous):

wouldnt blue be 0/10

OpenStudy (anonymous):

yeah it would.